# Bridging constrained random-phase approximation and linear response theory for computing Hubbard parameters

### v1

@@ Abstract
The predictive accuracy of popular extensions to density-functional theory (DFT) such as DFT+$U$ and DFT plus dynamical mean-field theory (DFT+DMFT) hinges on using realistic values for the screened Coulomb interaction $U$. Here, we present a systematic comparison of the two most widely used approaches to compute this parameter, linear response theory (LRT) and the constrained random-phase approximation (cRPA), using a unified framework based on the use of maximally localized Wannier functions to define the underlying basis sets. We demonstrate good quantitative agreement between LRT and cRPA in cases where the strongly interacting subspace corresponds to an isolated set of bands. Differences can be assigned to neglecting the response of the exchange-correlation potential in cRPA and the presence of additional screening channels in LRT. Moreover, we show that in cases with strong hybridization between interacting and screening subspaces, the application of cRPA becomes ambiguous and can lead to unrealistically small $U$ values, while LRT remains well-behaved. Our work clarifies the relation between both methods, sheds light on their strengths and limitations, and emphasizes the importance of using a consistent set of Wannier orbitals to ensure transferability of $U$ values between different implementations.
@@
An accurate description of many quantum materials [1] relies on correctly capturing the strong electron-electron interaction among a subset of electrons inhabiting the states around the Fermi level. The method of choice is to divide the electrons into strongly and weakly interacting subspaces, where the latter can be treated within common approximations to density-functional theory (DFT) [2, 3], while the strongly interacting subspace is augmented with a more accurate treatment of the local electron-electron interaction, resulting in the widely used DFT+$U$ [4, 5, 6, 7] or DFT plus dynamical mean-field theory (DFT+DMFT) methods [8, 9, 10].
@@
The accuracy and predictive capabilities of these techniques depend on a realistic estimation of the screened Coulomb interaction within the strongly interacting subspace, which is typically parametrized in terms of an on-site Hubbard $U$ (and potentially an on-site Hund’s coupling $J$ [5] and/or an inter-site interaction [11]). Consequently, several first-principles approaches to compute Hubbard parameters have been developed, including constrained DFT [12, 13, 14, 15, 16, 17, 18, 19, 20], linear response theory (LRT) [21, 22, 23, 24, 25, 26], Hartree-Fock-based methods [27, 28, 29, 30, 31, 32], and the constrained random-phase approximation (cRPA) [33, 34, 35, 36]. In particular, LRT and cRPA are widely used in practical applications. Notably, LRT is used almost exclusively for DFT+$U$ calculations, in particular for high-throughput applications of DFT+$U$ [37, 7, 38, 39], while cRPA is utilized more often in the context of DFT+DMFT calculations [40, 41, 42].
@@
This raises the question to what extent LRT and cRPA can be viewed as equivalent. While some attempts were made to compare different methods for computing Hubbard parameters [36, 43, 44], there is still lack of understanding and consensus regarding the similarities and differences between them. So far, an explicit quantitative comparison of the calculated $U$ parameters has been hindered by the fact that specific implementations generally use different basis orbitals to represent the strongly interacting subspace, with the $U$ parameter depending very sensitively on the specific choice of these orbitals [7, 45, 37, 46]. Many DFT codes implementing the LRT method use internally defined localized orbitals (e.g., orthogonalized atomic orbitals [45] or partial wave projectors [47]), while cRPA calculations are often based on Wannier functions [48, 49, 50, 51].
@@
In this letter, we systematically compare $U$ values obtained from LRT and cRPA. We specifically address the need for a consistent definition of the interacting subspace and a consistent choice of Hubbard projectors, which represent the corresponding localized basis orbitals. To this end, we extend the LRT calculation of the $U$ parameter using density-functional perturbation theory [22, 24] within Quantum ESPRESSO  [52, 53, 54] by interfacing it with Wannier90 [55] in order to employ maximally localized Wannier functions (MLWFs) [56] as Hubbard projectors. This enables a direct quantitative comparison between LRT and cRPA, the latter being evaluated using Quantum ESPRESSO in combination with RESPACK [57, 58]. We demonstrate that, for cases where the division of the interacting and screening subspaces corresponds to isolated or nearly-isolated sets of bands, LRT and cRPA yield remarkably similar results. However, we also show that the two approaches are not strictly equivalent and identify the corresponding differences. For cases where the two subspaces hybridize strongly, the application of cRPA becomes problematic, leading to unrealistically small $U$ values, whereas LRT remains largely unaffected.
@@
We begin by briefly summarizing the basic terminology of both approaches. An essential quantity in both methods is the charge susceptibility, or response tensor, which characterizes how a system of interacting electrons reacts to external perturbations. Thereby, one distinguishes between the bare susceptibility, $\chi_{0}$, which does not consider the change in the electronic interactions caused by the charge rearrangement, and the full susceptibility, $\chi$, which considers both the external perturbation and the effect of the resulting electron redistribution.
@@
In LRT, these susceptibilities are obtained by directly calculating the response of the Kohn-Sham (KS) system to small perturbations, with $\tilde{\chi}^{IJ}=\partial n^{I}/\partial\lambda^{J}$ defined as change in occupation of the Hubbard projector orbitals on site $I$ due to a local potential shift on site $J$, $\lambda^{J}$, acting only within the interacting subspace. The Hubbard parameter for site $I$ is then defined as the corresponding diagonal element of the difference of the inverse susceptibility matrices [7, 21]:
@@
where $\tilde{\chi}_{0}$ is calculated without considering the change in the KS potential resulting from the charge rearrangement. Note that here the tilde indicates that the corresponding susceptibilities only describe the “coarse grained” density response on a per site basis.
The cRPA formalism [35, 36] emphasizes the division of the electronic Hilbert space into the interacting subspace, $\mathcal{D}$, (also called “target” subspace or “correlated” subspace in the context of DFT+DMFT) and the screening subspace (also called “rest” subspace), mirroring the corresponding division within DFT+$U$ and DFT+DMFT.  [^1] In some cases, a different subdivision of the Hilbert space has been used to define the screening subspace within cRPA and to define the interacting subspace orbitals [50]. Here, we always use an identical subdivision to define both screening and interacting subspace. Within RPA, the bare susceptibility can be written as a sum over all possible particle-hole excitations in an independent particle basis:
where $|\psi_{n}\rangle$, $\epsilon_{n}$, and $f_{n}$ are the KS states, their energies, and occupations, respectively. One can split the bare susceptibility into two parts, $\chi_{0}=\chi_{0}^{\mathcal{D}}+\chi_{0}^{\mathcal{R}}$, where $\chi_{0}^{\mathcal{D}}$ contains only particle-hole excitations within the interacting subspace, and $\chi_{0}^{\mathcal{R}}$ contains all other transitions. The partially screened Coulomb interaction acting within the $\mathcal{D}$ subspace is then obtained as $W^{\mathcal{R}}=\left[1-V\chi_{0}^{\mathcal{R}}\right]^{-1}V$, where $V=1/|\mathbf{r}-\mathbf{r^{\prime}}|$ is the unscreened Coulomb interaction. Static interaction parameters are computed as matrix elements of $W^{\mathcal{R}}(\omega=0)$ with the Hubbard projector functions $\phi^{I}_{i}(\mathbf{r})$, i.e., the specific basis orbitals used to represent $\mathcal{D}$, corresponding to orbital $i$ on site $I$ [35]:
×$\times W^{\mathcal{R}}(\omega=0,\mathbf{r},\mathbf{r^{\prime}})\phi_{k}^{K}({\bf{r}}^{\prime})^{*}\phi_{l}^{L}({\bf{r}}^{\prime})\ d{\bf{r}}\ d{\bf{r}}^{\prime}.$ (3)
For practical calculations, one usually defines the Hubbard $U$ and the Hund’s exchange parameter $J$ by suitably averaging over matrix elements belonging to a single site. Here we use the Slater parametrization, which is often employed in DFT+$U$ calculations using the full $d$ shell [5], where $U$ is defined as average over both inter- and intra-orbital matrix elements (see, e.g.[59]):
$U_{\text{cRPA}}^{I}$ $=\frac{1}{M^{2}}\sum_{ij}(W^{\mathcal{R}})^{IIII}_{iijj},\quad\quad$ (4)
$U^{I}_{\text{cRPA}}-J^{I}_{\text{cRPA}}$ $=\frac{1}{M(M-1)}\sum_{ij}\left[(W^{\mathcal{R}})^{IIII}_{iijj}-(W^{\mathcal{R}})^{IIII}_{ijji}\right].$ (5)
Here, $M$ is the number of interacting orbitals on site $I$. We note that DFT+DMFT calculations sometimes employ the so-called Kanamori parametrization, where $U$ is defined as average over only the inter-orbital matrix elements, i.e., as $\tfrac{1}{M}\sum_{i}(W^{\mathcal{R}})^{IIII}_{iiii}$, which is larger than $U^{I}_{\text{cRPA}}$ defined in Eq. (4) by about $\tfrac{8}{7}J^{I}_{\text{cRPA}}$ [42]. On the other hand, LRT is typically discussed in the context of a further simplified form of the interaction tensor introduced by Dudarev et al. [6], with a single effective Hubbard parameter which can be identified as $U_{\text{eff}}=U-J$, with $U$ and $J$ corresponding to the Slater parametrization. Accordingly, in this work, we will compare $U_{\text{LRT}}$ with both $U_{\text{cRPA}}$ as well as with $U_{\text{cRPA}}-J_{\text{cRPA}}$.
@@
We focus on two different materials: KCuF3 and Sr2FeO4. KCuF3 has been studied using both DFT+$U$ and DFT+DMFT [5, 60], due to its Jahn-Teller-distorted structure and associated orbital order. For the purpose of this work, we consider KCuF3 in the high symmetry cubic perovskite structure [shown in Fig. 1(a)], both for ease of computation and because it exhibits the prototypical band structure of a transition metal (TM) perovskite, with a clear separation between bands with dominant Cu $d$ and F $p$ character [see Fig. 1(b)]. Sr2FeO4 crystallizes in a layered perovskite structure [shown in Fig. 1(a)] [61, 62], and is of interest due to its structural and electronic similarity to the unconventional superconductor Sr2RuO4 [42]. The corresponding KS bands [see Fig. 1(c)] exhibit a small overlap between bands with dominant Fe $d$ and dominant O $p$ character. As reported recently, the computed value of $U_{\text{cRPA}}$ for the Fe $d$ states is too small to reproduce the experimentally observed insulating state of Sr2FeO4 within DFT+DMFT [42].
@@
Figure 2: Interaction parameters obtained from LRT and cRPA for (a) KCuF3 amd (b) Sr2FeO4 using different choices for the Hubbard projectors (frontier, localized $d$-only and $d$-$p$). Blue, orange, and green columns corresponds to $U_{\text{LRT}}$, $U_{\text{cRPA}}-J_{\text{cRPA}}$, and $U_{\text{cRPA}}$, respectively. For Sr2FeO4, Oxy refers to the oxygen within the Fe planes, while Oz refers to the apical oxygen.
@@
In Fig. 2, we summarize the calculated values for $U_{\text{LRT}}$, $U_{\text{cRPA}}$, and $U_{\text{cRPA}}-J_{\text{cRPA}}$ for both KCuF3 (a) and Sr2FeO4 (b) and the three different choices for the interacting subspace. We first focus on the localized $d$-$p$ and the frontier orbital basis, which completely span the KS bands within the respective energy windows [see Fig. 1 (b) and (c)]. The interaction parameters obtained with LRT (shown in blue) are quantitatively rather similar to those obtained with cRPA. This holds for both basis sets and for both materials. Best agreement is typically observed between $U_{\text{LRT}}$ and $U_{\text{cRPA}}-J_{\text{cRPA}}$, which supports the notion that $U_{\text{LRT}}$ should be viewed as effective $U-J$. The agreement is particularly good for the localized $d$-$p$ basis, for which $U_{\text{LRT}}$ and $U_{\text{cRPA}}-J_{\text{cRPA}}$ differ by less than $3\%$, except for the F site in KCuF3. For the frontier case the difference is somewhat larger, around $12\%$. The large difference between $U_{\text{LRT}}$ and $U_{\text{cRPA}}-J_{\text{cRPA}}$ for the F $p$ states in KCuF3 in the $d$-$p$ basis, can be explained by the known tendency of LRT to overestimate $U$ in cases with essentially completely full (or completely empty) shells, where the applicability of the linear response approach becomes problematic [43, 63, 7]. Indeed, the occupation of the F $p$ states in KCuF3 is nearly 0.99 electrons per orbital in our $d$-$p$ basis, while in Sr2FeO4, the stronger hybridization with the TM $d$ states reduces the average occupation of the O $p$ states to $\sim$ 0.91, such that the applicability of LRT is unproblematic.
To better understand the good quantitative agreement as well as the remaining differences between LRT and cRPA for the frontier and localized $d$-$p$ cases, one can introduce a more general response tensor describing the full orbital-resolved response to a general non-local orbital-dependent perturbation [7]. As pointed out in [7], the LRT susceptibility can be viewed as a coarse-grained version of this generalized response tensor. One can further show (see the supplementary material [64] for full details), that for cases where the interacting subspace can be identified with a completely isolated set of bands, the bare cRPA susceptibility, $\chi_{0}^{\mathcal{R}}$, becomes equivalent to the generalized bare susceptibility restricted to the interacting subspace. If one then neglects the exchange-correlation contribution to the full susceptibility, which is included in LRT and in the generalized response function but is neglected in cRPA, $U_{\text{LRT}}$ can be related to the cRPA interaction tensor from Eq. (3) through the expression:
$U_{\text{LRT}}^{I}=\sum^{ZT}_{zt}\sum^{OPQS}_{opqs}\\ (\tilde{\chi}_{0}^{-1})^{IZ}\left[({\chi_{0}})_{zzpo}^{ZZPO}(W^{\mathcal{R}})_{opqs}^{OPQS}{\chi}_{sqtt}^{SQTT}\right](\tilde{\chi}^{-1})^{TI}\ .$ (6)
Here, $\chi_{0}$ and $\chi$ refer to the generalized susceptibilities, expressed in the orbital basis corresponding to the interacting subspace, while $\tilde{\chi}_{0}$ and $\tilde{\chi}$ refer to the coarse-grained LRT susceptibility matrices, with $\tilde{\chi}^{IR}=\sum_{ir}\chi^{IIRR}_{iirr}$ [7]. Rearranging the terms in Eq. (6), allows to write $U^{I}_{\mathrm{LRT}}=U^{I}_{\mathrm{cRPA}}+\Delta U^{I}$, where $U_{\text{cRPA}}$ arises from isolating terms containing only bi-diagonal components of the generalized susceptibility tensor (i.e., elements of the form $\chi_{sstt}^{SSTT}$) and replacing $(W^{\mathcal{R}})^{PPSS}_{ppss}$ with orbital averages defined analogous to Eq. (4) (see [64]). $\Delta U^{I}$ contains all other contributions arising from elements of the generalized susceptibility tensor that are not purely bi-diagonal. These include local exchange terms containing matrix elements of the form $(W^{\mathcal{R}})^{OOOO}_{oppo}$, as well as additional on-site and inter-site off-diagonal terms associated with particle-hole excitations within the interacting subspace that are excluded within cRPA.
To summarize, for the case of an isolated interacting subspace, and if the exchange correlation contribution to the response is neglected, $U_{\text{LRT}}$ is equal to $U_{\text{cRPA}}$ plus all the additional terms included in $\Delta U$. The overall good quantitative agreement between LRT and cRPA for the frontier and $d$-$p$ subspaces indicate that, at least for the materials considered here, these differences are small, in particular if the local exchange terms in Eq. (6) are taken into account in an effective way by considering $U_{\text{cRPA}}-J_{\text{cRPA}}$. However, we also note that test calculations where we removed the exchange-correlation part of the response within LRT led to an increase of $U_{\text{LRT}}$ by about 5-10%, depending on the specific case. This shows that both this effect as well as the additional terms in Eq. (6) can contribute notably to the difference between $U_{\text{LRT}}$ and $U_{\text{cRPA}}$.
Lastly, we discuss the localized $d$-only case (second set of columns in both panels of Fig. 2), for which LRT and cRPA lead to drastically different results. Both $U_{\text{cRPA}}$ and $U_{\text{cRPA}}-J_{\text{cRPA}}$ are very small in this case, with the latter even taking negative values for Sr2FeO4. In contrast, $U_{\text{LRT}}$ is much larger and comparable to the values obtained for the localized $d$-$p$ basis.

### v2

@@ Abstract
The predictive accuracy of popular extensions to density-functional theory (DFT) such as DFT+$U$ and DFT plus dynamical mean-field theory (DFT+DMFT) hinges on using realistic values for the screened Coulomb interaction $U$. Here, we present a systematic comparison of the two most widely used approaches to compute this parameter, i.e. linear response theory (LRT) and the constrained random-phase approximation (cRPA), using a unified framework based on the use of maximally localized Wannier functions. We show that the $U$ in LRT and cRPA can differ as much as 30%. We demonstrate that this discrepancy arises from two main differences: neglecting the response of the exchange-correlation potential in cRPA and additional excitation channels in LRT. By taking these differences into account, we can achieve near perfect agreement between the two techniques. Moreover, we show that in cases with strong hybridization between interacting and screening subspaces, the application of cRPA becomes ambiguous and can lead to unrealistically small $U$ values, while LRT remains well-behaved. Our work formally connects both methods, sheds light on their strengths and limitations, and emphasizes the importance of using a consistent set of Wannier orbitals to ensure transferability of $U$ values between different implementations.
@@
An accurate description of many quantum materials [Keimer/Moore:2017] relies on correctly capturing the strong electron-electron interaction among a subset of electrons inhabiting the states around the Fermi level. The method of choice is to divide the electrons into strongly and weakly interacting subspaces, where the latter can be treated within common approximations to density-functional theory (DFT) [Hohenberg:1964, Kohn:1965], while the strongly interacting subspace is augmented with a more accurate treatment of the local electron-electron interaction, resulting in the widely used DFT+$U$ [anisimov_band_1991, Liechtenstein:1995, dudarev_electronenergyloss_1998a, himmetoglu_hubbardcorrected_2013] or DFT plus dynamical mean-field theory (DFT+DMFT) methods [Georges:1996, kotliar_electronic_2006, held_electronic_2007].
@@
The accuracy and predictive capabilities of these techniques depend on a realistic estimation of the screened Coulomb interaction within the strongly interacting subspace, which is typically parametrized in terms of an on-site Hubbard $U$. Consequently, several first-principles approaches to computing Hubbard parameters have been developed, including constrained DFT [Dederichs:1984, Mcmahan:1988, Gunnarsson:1989, Hybertsen:1989, Gunnarsson:1990, Pickett:1998, Solovyev:2005, Nakamura:2006, Shishkin:2016], linear response theory (LRT) [cococcioni_linear_2005, timrov_hubbard_2018, linscott_role_2018, timrov_selfconsistent_2021, timrov_hp_2022, Macke:2024], Hartree-Fock-based methods [Mosey:2007, Mosey:2008, Andriotis:2010, Agapito:2015, TancogneDejean:2020, Lee:2020], and the constrained random-phase approximation (cRPA) [Springer:1998, Kotani:2000, aryasetiawan_frequencydependent_2004, aryasetiawan_calculations_2006, Scott2024]. In particular, LRT and cRPA are widely used in practical applications. Notably, LRT is used almost exclusively for DFT+$U$ calculations, in particular for high-throughput applications of DFT+$U$ [kirchner-hall_extensive_2021, himmetoglu_hubbardcorrected_2013, capdevila-cortada_performance_2016, moore_highthroughput_2024], with few exceptions [Ghosh2023], while cRPA is utilized more often in the context of DFT+DMFT calculations [seth_renormalization_2017, souto-casares_oxygen_2021, kazemi-moridani_strontium_2024a].
@@
Crucially, it has not been established to what extent LRT and cRPA can be viewed as equivalent.While some attempts were made to compare different methods for computing Hubbard parameters [aryasetiawan_calculations_2006, tesch_hubbard_2022, Liu:2023, Yang_2025_arxiv], there is still lack of understanding and consensus regarding the similarities and differences between them. So far, an explicit quantitative comparison of the calculated $U$ parameters has been hindered by the fact that specific implementations generally use different basis orbitals to represent the strongly interacting subspace, with the $U$ parameter depending very sensitively on the specific choice of these orbitals [himmetoglu_hubbardcorrected_2013, Carta_et_al:2025, kirchner-hall_extensive_2021, Mahajan:2021]. Many DFT codes implementing the LRT method use internally defined localized orbitals (e.g., orthogonalized atomic orbitals [Carta_et_al:2025] or partial wave projectors [Bengone_et_al:2000]), while cRPA calculations are often based on Wannier functions [miyake_initio_2009, sakuma_firstprinciples_2013, miyake_forbital_2008, miyake_screened_2008].
@@
In this letter, we systematically compare $U$ values obtained from LRT and cRPA. We specifically address the need for a consistent definition of the interacting subspace and a consistent choice of Hubbard projectors, which represent the corresponding localized basis orbitals. To this end, we extend the LRT calculation of the $U$ parameter using density-functional perturbation theory [timrov_hubbard_2018, timrov_selfconsistent_2021] within Quantum ESPRESSO  [giannozzi_quantum_2009a, giannozzi_advanced_2017a, Giannozzi:2020] by interfacing it with Wannier90 [pizzi_wannier90_2020] in order to employ maximally localized Wannier functions (MLWFs) [Marzari:2012] as Hubbard projectors. This enables a direct quantitative comparison between LRT and cRPA, the latter being evaluated using Quantum ESPRESSO in combination with RESPACK [nakamura_respack_2021a, kurita_interface_2023]. We first show mathematically that the definitions of $U$ in the two approaches are not strictly equivalent, but that, for cases where the interacting and screening subspaces correspond to isolated or nearly-isolated sets of bands, one can clearly identify the corresponding differences. We then demonstrate numerically, that LRT and cRPA can be brought in remarkable quantitative agreement, if these differences are properly accounted for. For cases where the two subspaces hybridize strongly, the application of cRPA is problematic, leading to unrealistically small $U$ values, whereas LRT remains largely unaffected.
@@
An essential quantity in both methods is the charge susceptibility, or response tensor, which characterizes how a system of interacting electrons reacts to external perturbations. Thereby, one distinguishes between the bare susceptibility, $\chi_{0}$, which does not consider the change in the electronic interactions caused by the charge rearrangement, and the full susceptibility, $\chi$, which considers both the external perturbation and the effect of the resulting electron redistribution.
Within RPA, the bare susceptibility can be written as a sum over all possible electron-hole excitations in an independent particle basis:
where $|\psi_{n}\rangle$, $\epsilon_{n}$, and $f_{n}$ are the Kohn-Sham (KS) states, their energies, and occupations, respectively. The cRPA formalism [aryasetiawan_frequencydependent_2004, aryasetiawan_calculations_2006] then emphasizes the division of the electronic Hilbert space into the interacting subspace, $\mathcal{D}$, and the screening subspace, mirroring the corresponding division within DFT+$U$ and DFT+DMFT.  [^1] In some cases, a different subdivision of the Hilbert space has been used to define the screening subspace within cRPA and to define the interacting subspace orbitals [miyake_forbital_2008]. Here, we always use an identical subdivision to define both screening and interacting subspace. The bare susceptibility is then split into two parts, $\chi_{0}=\chi_{0}^{\mathcal{D}}+\chi_{0}^{\mathcal{R}}$, where $\chi_{0}^{\mathcal{D}}$ contains only electron-hole excitations within the interacting subspace, and $\chi_{0}^{\mathcal{R}}$ contains all other transitions. The partially screened Coulomb interaction acting within the $\mathcal{D}$ subspace is then obtained as $W^{\mathcal{R}}=\left[1-V\chi_{0}^{\mathcal{R}}\right]^{-1}V$, where $V=1/|\mathbf{r}-\mathbf{r^{\prime}}|$ is the unscreened Coulomb interaction. For the following it is also useful to define the full susceptibility within the interacting subspace, $\chi^{\mathcal{D}}$, through the Dyson equation $W^{\mathcal{R}}=(\chi_{0}^{\mathcal{D}})^{-1}-(\chi^{\mathcal{D}})^{-1}$.
Static interaction parameters are computed as matrix elements of $W^{\mathcal{R}}(\omega=0)$ with the Hubbard projector functions $\phi^{I}_{i}(\mathbf{r})$, i.e., the specific basis orbitals used to represent $\mathcal{D}$, corresponding to orbital $i$ on site $I$ [aryasetiawan_frequencydependent_2004]:
$\times W^{\mathcal{R}}(\omega=0,\mathbf{r},\mathbf{r^{\prime}})\phi_{k}^{K}({\bf{r}}^{\prime})^{*}\phi_{l}^{L}({\bf{r}}^{\prime})\ d{\bf{r}}\ d{\bf{r}}^{\prime}.$ (2)
Within the Slater parametrization of the local interaction, which is often employed in calculations using the full $d$ shell [Liechtenstein:1995], $U$ is defined as average over both inter- and intra-orbital matrix elements (see, e.g., Ref. [merkel_calculation_2024a]):
$U_{\text{cRPA}}^{I}$ $=\frac{1}{(N^{I})^{2}}\sum_{ij}(W^{\mathcal{R}})^{IIII}_{iijj}\ ,$ (3)
where $N^{I}$ is the number of interacting orbitals on site $I$.
@@
On the other hand, in LRT, the relevant susceptibilities are obtained by directly calculating the response of the KS system to small localized perturbations, with $\tilde{\chi}^{IJ}=\partial n^{I}/\partial\lambda^{J}$ defined in terms of the change in occupation of the Hubbard projector orbitals on site $I$ due to a local potential shift on site $J$, $\lambda^{J}$, acting only within the interacting subspace. The Hubbard parameter for site $I$ is then defined as the corresponding diagonal element of the difference of the inverse susceptibility matrices [himmetoglu_hubbardcorrected_2013, cococcioni_linear_2005]:
@@
where $\tilde{\chi}_{0}$ is calculated without considering the change in the KS potential resulting from the charge rearrangement. Note that $\tilde{\chi}_{0}$ and $\tilde{\chi}$ describe the density-density response only on a per site basis, without any orbital dependence, i.e., they describe a purely monopolar response. However, as shown in Ref. [himmetoglu_hubbardcorrected_2013], these “coarse-grained” susceptibilities can be related to a more general, site- and orbital-dependent, 4-index response function. In Sec. I.F of the supplemental information (SI) [SI] we show that for cases where the interacting subspace $\mathcal{D}$ is formed from an isolated set of bands, the cRPA susceptibilities $\chi_{0}^{\mathcal{D}}$ and $\chi^{\mathcal{D}}$ can be identified with this general 4-index response function if one neglects the exchange-correlation (xc) contribution to the susceptibility (the latter is included in LRT but typically neglected in cRPA). The coarse grained response can then be written in terms of the orbital averages of the bi-diagonal components, $\tilde{\chi}_{0}^{IJ}=\sum_{ij}(\chi_{0}^{\mathcal{D}})_{iijj}^{IIJJ}$ and $\tilde{\chi}^{IJ}=\sum_{ij}(\chi^{\mathcal{D}})_{iijj}^{IIJJ}$ [SI].
@@
$U_{\text{LRT}}$ can then be related to the cRPA interaction tensor, Eq. (2), through:
$U_{\text{LRT}}^{I}=\sum^{ZT}_{zt}\sum^{OPQS}_{opqs}(\tilde{\chi}_{0}^{-1})^{IZ}\\ \times\left[({\chi^{\mathcal{D}}_{0}})_{zzpo}^{ZZPO}(W^{\mathcal{R}})_{opqs}^{OPQS}({\chi^{\mathcal{D}}})_{sqtt}^{SQTT}\right](\tilde{\chi}^{-1})^{TI}\ ,$ (5)
where all indices correspond to $\mathcal{D}$.
If we separate the cRPA susceptibilities into their locally “isotropic” component (i.e., the purely monopolar part that gives rise to $\tilde{\chi}_{0}$ and $\tilde{\chi}$) and a locally “non-isotropic” part, see Eq. (42) in the SI [SI], the isotropic contributions cancel with the corresponding coarse-grained susceptibilities in Eq. (5), resulting in the orbitally-averaged $U_{\mathrm{cRPA}}^{I}$ from Eq. (3) plus all remaining terms containing non-isotropic parts. Thus, we can write $U_{\text{LRT}}^{I}=U_{\text{cRPA}}^{I}+\Delta U^{I}$, where $\Delta U^{I}$ contains all non-isotropic contributions. These additional terms correspond to excitations completely inside the interacting subspace, and are thus excluded from the screening within cRPA. They represent processes that alter the “shape” of the charge density on an atomic site (orbital reorganizations, bond fluctuations) and also include local exchange terms containing matrix elements of the form $(W^{\mathcal{R}})^{OOOO}_{oppo}$.
Thus, in the case of well separated interacting and screening subspaces, we can identify two differences in the definition of the $U$ parameter between LRT and cRPA: (1) the xc contribution to the response is typically neglected in cRPA but included in LRT, and (2) the coarse-grained nature of LRT introduces additional terms $\Delta U^{I}$ related to non-isotropic excitations within $\mathcal{D}$.
We now investigate how these differences manifest numerically within realistic calculations, using two different materials as benchmarks: KCuF3 and Sr2FeO4. Sec. V of the SI [SI] contains results for two additional materials: NiO and CrO2.
KCuF3 has been studied using both DFT+$U$ and DFT+DMFT [Liechtenstein:1995, Pavarini/Koch/Lichtenstein:2008], due to its Jahn-Teller-distorted structure and associated orbital order. For the purpose of this work, we consider KCuF3 in the high symmetry cubic perovskite structure [shown in Fig. 1(a)], both for ease of computation and since it exhibits a prototypical band structure of a transition metal (TM) perovskite, with a clear separation between bands with dominant Cu $d$ and F $p$ character [see Fig. 1(b)]. Sr2FeO4 crystallizes in a layered perovskite structure [shown in Fig. 1(a)] [dann_structure_1993, dann_synthesis_1991], and is of interest due to its structural and electronic similarity to the unconventional superconductor Sr2RuO4 [kazemi-moridani_strontium_2024a]. The corresponding KS bands [see Fig. 1(c)] exhibit a small overlap between bands with dominant Fe $d$ and dominant O $p$ character. As reported recently, the computed value of $U_{\text{cRPA}}$ for the Fe $d$ states is too small to reproduce the experimentally observed insulating state of Sr2FeO4 within DFT+DMFT [kazemi-moridani_strontium_2024a].
Figure 2: Interaction parameters obtained from LRT and cRPA for (a) KCuF3 and (b) Sr2FeO4 using different choices for the Hubbard projectors (frontier, $d$-$p$, and $d$-only). Dark blue and dark red bars correspond to interaction parameters computed from a standard LRT and cRPA calculation respectively. The light blue bar corresponds to a modified LRT calculation, i.e. by removing the contribution of the xc kernel, while the light red bar corresponds to the modified cRPA calculation, i.e. obtained by using the coarse graining. For Sr2FeO4, Oxy refers to the oxygen within the Fe planes, while Oz refers to the apical oxygen. The coarse-grained cRPA results for the $d$-only model are not physically meaningful and are thus not included here.
@@
In Fig. 2, we summarize the calculated values for various definitions of $U$ for both KCuF3 (a) and Sr2FeO4 (b) and the three different choices for the interacting subspace. We first focus on the localized $d$-$p$ and the frontier orbital basis, which completely span the KS bands within the respective energy windows [see Fig. 1 (b) and (c)]. The interaction parameters obtained from LRT (shown in dark blue) are of comparable magnitude to those obtained from cRPA (in dark red), with $U_{\text{cRPA}}$ being larger than $U_{\text{LRT}}$ by around 10-30%. The only exception are the F $p$ states in KCuF3 in the $d$-$p$ basis, for which $U_{\text{LRT}}$ is about 7% larger. This last result can be explained by the known tendency of LRT to overestimate $U$ in cases with essentially completely full (or completely empty) shells, where the applicability of the linear response approach becomes problematic [tesch_hubbard_2022, yu_communication_2014, himmetoglu_hubbardcorrected_2013]. Indeed, the occupation of the F $p$ states in KCuF3 is nearly 1.0 electrons per orbital in our $d$-$p$ basis, while in Sr2FeO4, the stronger hybridization with the TM $d$ states reduces the average occupation of the O $p$ states to $\sim$ 0.9, such that the applicability of LRT is not problematic.
To relate these quantitative differences between $U_{\text{LRT}}$ and $U_{\text{cRPA}}$ to our mathematical considerations, we first analyze the effect of removing the response of the xc potential from the LRT calculation (light blue bars in Fig. 2), as explained in Sec. II in the SI [SI]. This generally increases $U_{\text{LRT}}$ by 55-$10\%$, thus bringing it closer to the cRPA value. To analyze the remaining difference to $U_{\text{cRPA}}$, we then perform a coarse-graining of the cRPA results by computing $\tilde{\chi}_{0}$ and $\tilde{\chi}$ from the 4-index tensors $\chi^{\mathcal{D}}_{0}$ and $\chi^{\mathcal{D}}$ (see Sec. I in the SI [SI]), and extract a corresponding $U$ via Eq. (4). The resulting value (light red bars in Fig. 2) is mathematically the same quantity as $U_{\text{LRT}}$ with the xc response neglected. Indeed, the corresponding values show very good quantitative agreement. They are nearly identical for KCuF3, while for Sr2FeO4 they exhibit residual discrepancies of about $5\%$. These small numerical differences can be assigned to the finite convergence threshold used in the calculations and the fact that RESPACK computes, in addition to all density-density terms, $(W^{\mathcal{R}})^{IIJJ}_{iijj}$ across all cells, only the exchange integrals, $(W^{\mathcal{R}})^{IJJI}_{ijji}$, in the first unit cell. Other elements are thus missing from our coarse-graining procedure. Overall, the remarkable quantitative agreement between modified LRT and cRPA for both the frontier and $d$-$p$ bases, despite using completely different implementations, strongly supports our analytical result.